### v1

@@ Abstract
Dynamic treatment regimes have been proposed to personalize treatment decisions by utilizing historical patient data, but optimization can only be done over information available in the database. In contrast, the standard of care or physicians’ decisions may be complex algorithms based on information that is not available to researchers. It is thus meaningful to integrate the standard of care into the evaluation of treatment strategies, and previous works have suggested doing so through the concept of clinical utility. Here we will focus on the comparative component of clinical utility as the average outcome had the full population received treatment based on the proposed dynamic treatment regime in comparison to the full population receiving the “standard” treatment assignment mechanism, such as a physician’s choice. Clinical trials to evaluate clinical utility are rarely conducted, and thus, previous works have proposed an emulated clinical trial framework using observational data. However, only one simple estimator was previously suggested, and the practical details of how one would conduct this emulated trial were not detailed. Here, we illuminate these details and propose several estimators of clinical utility based on estimators proposed in the dynamic treatment regime literature. We illustrate the considerations and the estimators in a real data example investigating treatment rules for rheumatoid arthritis, where we highlight that in addition to the standard of care, the current medical guidelines should also be compared to any ‘optimal’ decision rule.
@@ 1 Introduction
Estimating the outcome of using covariates to guide treatment decisions has been explored by [4] and [5], among others, under the name of dynamic treatment regimes (DTRs). DTRs are rules for making a treatment decision based on a subset of past information. In this framework, one naturally wants to find the optimal treatment regime, which is the regime that maximizes or minimizes the outcome. In general, a DTR can involve time varying treatments but we limit our focus to the case of a treatment decision at a single time point. A simple example of a DTR in the single time point context is the guidelines of the [6] for rabies post-exposure prophylaxis that assigns treatment according to exposure covariates, such as the nature of the contact with the animal and the risk of rabies transmission. For lower-risk exposures (touching or feeding animals or licks on intact skin), no treatment is required. For moderate-risk exposures (minor scratches or abrasions), immediate vaccination is assigned. High-risk exposures (bites, scratches, or contact with broken skin or mucous membranes), require both vaccination and rabies immunoglobulin.
@@
Clinical utility is the average outcome had the full population received treatment based on a regime in comparison to the full population receiving the “standard” treatment assignment mechanism. This differs from the DTRs paradigm which does not generally include a comparison to the current standard of care. This can potentially be problematic, if, for example, the optimal treatment regime is not superior to the standard of care, and the standard of care is itself too complex to be considered as a regime under the optimization. Physician and patient’s choice can be a complex decision process involving information not collected as research data.
@@ 2 Preliminaries and Notation
Let $Y$ be the random variable for the outcome of interest, which is assumed to be binary (0/1) or continuous. Let $T$ be the discrete random variable for the observed treatment, with $k\geq 2$ levels. Let $\boldsymbol{Z}$ be a vector of measured covariates of dimension $d$. We will assume throughout that we have $n$ independent identically distributed sampled data points of $(T,\boldsymbol{Z},Y)$. We assume throughout that the observed data satisfies the causal graph represented by the directed acyclic graph below with the associated nonparametric structural equation models (NPSEM), with independent noise $U_{Z},U_{T}$ and $U_{Y}$.
@@
A DTR is a function $f$ that assigns a treatment $t$ based on a covariate vector $\boldsymbol{z}$. The potential outcome $Y(f(\boldsymbol{Z}))$ of an individual is the outcome if the individual is assigned the treatment that the DTR $f$ prescribes for the covariates $\boldsymbol{Z}$. After intervening with a DTR $f$ on the treatment, the equation of $T$ is replaced with $T:=f(\boldsymbol{Z})$ while the equations for $\boldsymbol{Z}$ and $Y$ stay the same. The DTR of the observed assignment mechanism is defined by $g_{T}$, which may be a complex unknown function of both measured and unmeasured covariates, defined as $U_{T}$ in the nonparametric structural equation. Thus, the expected outcome under the observed treatment assignment mechanism is $E[Y(g_{T}(\boldsymbol{Z},U))]=E(Y).$ There may also be clinical guidelines for patient treatment, which is also a DTR that we will denote $f_{cgl}$; we will denote the expected values of the outcome enforcing these guidelines as $E[Y(f_{cgl}(\boldsymbol{Z}))]$.
@@
This allows us to formally define the clinical utility as the difference in expected outcomes comparing two DTRs. In particular, the clinical utility of $f$ as defined in [3] is $E[Y(f(\boldsymbol{Z}))]-E[Y]$, this is in comparison to the observed assignment mechanism. It may be the case that the standard of care has changed over the course of the observation period, and thus the desired comparison is not to $E[Y]$. One may also wish to compare $f$ to the clinical guidelines $E[Y(f(\boldsymbol{Z}))]-E[Y(f_{cgl}(\boldsymbol{Z}))]$, which is also the clinical utility, as any new decision rule will at best become a new set of guidelines. We discuss these issues further in Section 4.4. In order to identify and estimate the clinical utility:
We make the following assumptions:
  1. 1.
    Positivity: For any observed $\boldsymbol{z}$ we have that $0<P(T=t\mid\boldsymbol{Z}=\boldsymbol{z})$ for all values of $t$ that we wish to consider.
  2. 2.
    No unmeasured confounders: For any treatment $t$, $T$ is independent of the potential outcome $Y(t)$ conditioned on the measured covariates $\boldsymbol{Z}$ which is evident from the DAG above. That is $\forall t,\quad T\perp Y(t)\mid\boldsymbol{Z}$.
@@
Under these assumptions and the NPSEM given above, we can identify $E[Y(f(\boldsymbol{Z}))]$ and estimate it using existing estimators.
@@ 2.1 Estimators
Let $f$ be a DTR. Then, under the assumption of positivity and no unmeasured confounding, the estimand of the potential outcome under $f$ can be written as
@@
See ([7] p. 59) or in the supplementary materials section 2 for reference. This gives rise to a consistent and asymptotically unbiased estimator, given by:
@@
For estimator 1, the propensity score model for $\hat{p}(T=t_{i}\mid\boldsymbol{z}_{i})$ has to be correctly specified. Similarly, for estimator 2 it is the binary propensity score model for $\hat{p}(\mathds{1}[t_{i}=f(\boldsymbol{z}_{i})]=1\mid\boldsymbol{z}_{i})$ that has to be correctly specified.
@@
In the first, the outcome model $h_{b}(\cdot)$, and in the second, $h_{nb}(\cdot)$ for $Y$, must be correctly specified. For $h_{b}$ the predictor variables include $\mathds{1}[T=f(\boldsymbol{z}_{i})]$, which we use to directly predict $\hat{Y}_{i}(\mathds{1}[T=f(\boldsymbol{z}_{i})]=1)$ for all subjects. In contrast, in $h_{nb}$, we include $T$ as a predictor variable, and predict $\hat{Y}_{i}(T=f(\boldsymbol{z}_{i}))$ for each subject. We leave the outcome models otherwise unspecified as $h_{b}(\cdot)$ and $h_{nb}(\cdot)$, which could be anything from logistic regression to random forest.
@@ 3 Simulation study
The covariate $\boldsymbol{Z}=(Z_{1},Z_{2})$ are sampled from a discrete uniform distribution $Z_{1}\sim\mathcal{U}(0,1)$ in discrete steps of $0.2$ and a binomial distribution $Z_{2}\sim\mathcal{B}(0.7)$. Next, we assign each person a treatment (which will represent the standard of care). This observed treatment allocation differs in Setting 1-3. For a comprehensive description of the data generation process, see supplementary material section 3. Last, we need to simulate the outcome. The generation of the outcome $Y$ relies on the risk function defined by:
$P(Y=1\mid\mathbb{z},t)=\begin{cases}0.5+0.5z_{1}-0.5z_{2}&\text{if }t=1\\ 0.65-0.5z_{2}&\text{ if }t=2\\ 1-0.5z_{1}-0.5z_{2}&\text{ if }t=3\\ \end{cases}$ (3)
@@
  • Setting 2 (S2): Contrasting, in this setting, the clinical utility is zero. Despite employing a different DTR, individual treatments are allocated in a manner that produces an identical counterfactual average outcome to that of the standard of care. This scenario presents a nullifying result.
@@
The case where the DTR under investigation performs worse than the SOC is analogous to Setting 1 and 3 but with a change in sign of the clinical utility.
@@ 3.1 Results
The results of our simulation study are displayed in Table 1. Notably, for a substantial sample size of $n=2000$, all estimators across Settings 1 to 3 exhibit minimal bias, coupled with a commendable coverage level hovering around $95\%$. However, as sample sizes decrease, we observe a gradual increase in the standard error. The performance degradation of the two IPW estimators is particularly noteworthy, becoming apparent at the smallest sample size of $n=200$, accompanied by a notable increase in bias and standard error, which consequently leads to a decrease in coverage accuracy. With a sample size of $n=200$, we observe that the variance of the $\text{gc}_{nb}$ estimator is larger than that of both IPW estimators, which is surprising given our expectation that the IPW estimators would degrade more quickly in general. In our case this can be explained by the $\text{gc}_{b}$ model’s inconsistency to fit all 44 parameters (due to all the interactions) at such small sample size. In contrast, the simpler $\text{ipw}_{n}$ model, which requires only 22 parameters, is reliably fitting all parameters. On the other hand, the $\text{gc}_{nb}$ estimator shows the smallest variance because its correctly specified outcome model is exactly the risk function 3 of the data generating mechanism, which makes it impossible to be outperformed. The estimators $ipw_{b}$ and $ipw_{n}b$ are almost identical in S1-S3. Since $h_{b}$ and $h_{nb}$ are the correctly specified outcome models of the propensity scores, they are the same for individuals we sum over, i.e. the ones for which $f(z)=t$. In the misspecified setting $ipw_{b}$ and $ipw_{nb}$ differ as their outcome models for the propensity scores are differently (mis)specified.
@@
Estimator B×$\times 10^{2}$ SE×$\times 10^{1}$ Co B×$\times 10^{2}$ SE×$\times 10^{1}$ Co B×$\times 10^{2}$ SE×$\times 10^{1}$ Co
S1 $ipw_{b}$0.01 0.14 0.95 0.65 0.29 0.91 2.80 0.45 0.78
$ipw_{nb}$0.01 0.14 0.95 0.65 0.29 0.91 2.80 0.45 0.77
$gc_{b}$0.02 0.14 0.94 0.05 0.29 0.97 0.31 0.56 0.99
@@ 4.1 Data
The outcome of interest is the log C-reactive protein (CRP) level measured at the earliest time point between 3 months after the treatment switch and up to 1 year after the treatment switch. CRP is a marker of inflammation, hence a smaller outcome is more desirable. Covariates which were thought to possibly influence the outcome and/or the treatment selection included in the models are the phase I treatment, last measured CRP before treatment switch, gender, indicator of being married, indicator of being an immigrant or child of an immigrant, years since RA diagnosis, history of cardiovascular disease, history of chronic obstructive pulmonary disease, and history of osteoperosis. A description of the sample is in Table 2.
Table 2: Descriptive statistics of the Danish register data by phase II observed treatment group.
@@ 4.2 Reasoning about time zero
The ideal trial that we would be emulating would have a clearly defined time point where subjects enter the trial and are randomized to either the estimated optimal DTR or the SOC. At the time of entry into the trial, they would also have the necessary biomarkers or other data measured that are inputs to the DTR or SOC regimes. Thus we need to have a similar time zero defined in our emulated trial; please see Hernan and Robins[14] and [15] [16]for more discussion of this point in general. The DTR setting differs in an important way from the standard emulated treatment trial because of the requirement of having the necessary biomarkers measured to apply the DTR. If the DTR requires high-dimensional data or a large number of biomarkers, it may be difficult to identify individuals in an observational dataset who have all the required biomarkers measured within a small time window.
@@ 4.3 Inclusion and exclusion criteria
We include adults who were diagnosed with RA before 1 January 2016 (ICD-10 codes DM05, DM06), and who started first-line treatment with hydroxycloraquine, methotrexate, or sulfasalazine, and who changed or added a biologic or csDMARD during the years 2016, 2017, and 2018. This treatment changing instance would correspond with the phase II treatment decision as described by the EULAR 2023 recommendations [9]. Since the guidelines are based on prognostic factors, we restricted to individuals with either or both rheumatoid factor and C-reactive protein levels measured within the 6 month time window before the treatment switch. We define the outcome as the biomarker CRP value after at least 3 month and no more than one year, after the switch of treatment.

### v2

@@ Abstract
Dynamic treatment regimes have been proposed to personalize treatment decisions by utilizing historical patient data, but they may not always improve on the current standard of care. It is thus meaningful to integrate the standard of care into the evaluation of treatment strategies, and previous works have suggested doing so through the concept of clinical utility. Here we will focus on the comparative component of clinical utility as the average outcome had the full population received treatment based on the proposed dynamic treatment regime in comparison to the full population receiving the “standard” treatment assignment mechanism, such as a physician’s choice. Clinical trials to evaluate clinical utility are rarely conducted, and thus, previous works have proposed an emulated clinical trial framework using observational data. However, only one simple estimator was previously suggested, and the practical details of how one would conduct this emulated trial were not detailed. Here, we illuminate these details and propose several estimators of clinical utility based on estimators proposed in the dynamic treatment regime literature. We illustrate the considerations and the estimators in a real data example investigating treatment rules for rheumatoid arthritis, where we highlight that in addition to the standard of care, the current medical guidelines should also be compared to any estimated “optimal” decision rule.
@@ 1 Introduction
Estimating the outcome of using covariates to guide treatment decisions has been explored by [4] and [5], among others, under the name of dynamic treatment regimes (DTRs). DTRs are rules for making a treatment decision based on a subset of past information. In this framework, one naturally wants to find the optimal treatment regime, which is the regime that maximizes or minimizes the outcome. In general, a DTR can involve time varying treatments but we limit our focus to the case of a treatment decision at a single time point. A simple example of a DTR in the single time point context is the guidelines for rabies post-exposure prophylaxis that assigns treatment according to exposure covariates, such as the nature of the contact with the animal and the risk of rabies transmission [6]. For lower-risk exposures (touching or feeding animals or licks on intact skin), no treatment is required. For moderate-risk exposures (minor scratches or abrasions), immediate vaccination is assigned. High-risk exposures (bites, scratches, or contact with broken skin or mucous membranes), require both vaccination and rabies immunoglobulin.
@@
Clinical utility is the average outcome had the full population received treatment based on a regime in comparison to the full population receiving the “standard” treatment assignment mechanism. This differs from the DTRs paradigm which does not generally include a comparison to the current standard of care. This can potentially be problematic, if, for example, the optimal treatment regime is not superior to the standard of care, and the standard of care is itself too complex to be considered as a regime under the optimization. Physician and patient’s choice may be a complex decision process involving information not collected as research data.
@@ 2 Preliminaries and Notation
Let $Y$ be the random variable for the outcome of interest, which is assumed to be binary (0/1) or continuous. Let $T$ be the discrete random variable for the observed treatment, with $k\geq 2$ levels. Let $\boldsymbol{Z}$ be a vector of measured covariates of dimension $d$. We will assume throughout that we have $n$ independent identically distributed sampled data points of $(T,\boldsymbol{Z},Y)$. We assume throughout that the observed data satisfies the causal graph represented by the directed acyclic graph below with the associated nonparametric structural equation models (NPSEMs), with independent noise $U_{Z},U_{T}$ and $U_{Y}$.
ZTY $Z$ $=g_{Z}(U_{z})$ $T$ $=g_{T}(\boldsymbol{Z},U_{T})$ $Y$ $=g_{Y}(\boldsymbol{Z},T,U_{Y}).$
The NPSEMs framework can be seen as a bridge between graphical and potential-outcome models [10]. By defining variables through arbitrary, and flexible, functional forms $(g_{Z},g_{T},g_{Y})$ rather than distributional or counterfactual assumptions alone, this approach explicitly characterizes the variables in the graph as intervenable quantities. This makes consistency — that the observed outcome equals the potential outcome under the observed treatment — an axiom rather than an assumption. Most of the remaining common counterfactual assumptions for identification of an estimand are then implied by the characteristics of the specific NPSEM. Here, mutually independent errors $(U_{Z},U_{T},U_{Y})$ and the lack of any common latent causes implies exchangeability, i.e. no unmeasured confounders. Positivity is still a needed assumption (defined below).
@@
A DTR is a function $f$ that assigns a treatment $t$ based on a covariate vector $\boldsymbol{z}$. The potential outcome $Y(f(\boldsymbol{Z}))$ of an individual is the outcome if the individual is assigned the treatment that the DTR $f$ prescribes for the covariates $\boldsymbol{Z}$. After intervening with a DTR $f$ on the treatment, the equation of $T$ is replaced with $T:=f(\boldsymbol{Z})$ while the equations for $\boldsymbol{Z}$ and $Y$ stay the same. The DTR of the observed assignment mechanism is defined by $g_{T}$, which may be a complex unknown function of both measured and unmeasured covariates. Thus, the expected outcome under the observed treatment assignment mechanism is $E[Y(g_{T}(\boldsymbol{Z},U))]=E(Y)$, which clearly does not fit the provided DAG and NPSEM if $U$ is associated with $Y$, but may fit the DAG if $U=U_{T}$. There may also be clinical guidelines for patient treatment, which is also a DTR that we will denote as $f_{cgl}$; we will denote the expected values of the outcome enforcing these guidelines as $E[Y(f_{cgl}(\boldsymbol{Z}))]$.
@@
This allows us to formally define the clinical utility as the difference in expected outcomes comparing two DTRs. In particular, the clinical utility of $f$ as defined in [3] is $E[Y(f(\boldsymbol{Z}))]-E[Y]$, this is in comparison to the observed assignment mechanism. It may be the case that the standard of care has changed over the course of the observation period, and thus the ideal comparison is not to $E[Y]$. One may also wish to compare $f$ to the clinical guidelines $E[Y(f(\boldsymbol{Z}))]-E[Y(f_{cgl}(\boldsymbol{Z}))]$, which is also the clinical utility, as any new decision rule will at best become a new set of guidelines. We discuss these issues further in Section 4.4.
@@
For identification and estimation the clinical utility, we assume Positivity, i.e. for any observed $\boldsymbol{z}$ we have that $0<P(T=t\mid\boldsymbol{Z}=\boldsymbol{z})$ for all values of $t$ that we wish to consider, in addition to the NPSEMs given above. Under these assumptions, we can identify $E[Y(f(\boldsymbol{Z}))]$ and estimate it using existing estimators.
@@ 2.1 Estimators
Let $f$ be a DTR. Then, under the assumption of positivity and the NPSEM above the estimand of the potential outcome under $f$ can be written as
@@
See ([7] p. 59) or in the Supplementary Materials Section 2 for details. This gives rise to a consistent and asymptotically unbiased estimator, given by:
@@
To obtain consistent estimates, the propensity score model for $\hat{p}(T=t_{i}\mid\boldsymbol{z}_{i})$ has to be correctly specified in 1 and, similarly, for estimator 2 it is the binary propensity score model for $\hat{p}(\mathds{1}[t_{i}=f(\boldsymbol{z}_{i})]=1\mid\boldsymbol{z}_{i})$ that has to be correctly specified.
@@
In the first, the outcome model $h_{b}(\cdot)$, and in the second, $h_{nb}(\cdot)$ for $Y$, must be correctly specified for consistent estimation. For $h_{b}$, the predictor variables include $\mathds{1}[T=f(\boldsymbol{z}_{i})]$, which we use to directly predict $\hat{Y}_{i}(\mathds{1}[T=f(\boldsymbol{z}_{i})]=1)$ for all subjects. In contrast, in $h_{nb}$, we include $T$ as a predictor variable, and predict $\hat{Y}_{i}(T=f(\boldsymbol{z}_{i}))$ for each subject. We leave the outcome models otherwise unspecified as $h_{b}(\cdot)$ and $h_{nb}(\cdot)$.
@@ 3 Simulation study
$\boldsymbol{z}=\begin{pmatrix}z_{1}\\ z_{2}\end{pmatrix}\in[0,1]\times\{0,1\}$ $\text{covariates}.$
The covariates $\boldsymbol{Z}=(Z_{1},Z_{2})$ are sampled from a discrete uniform distribution $Z_{1}\sim\mathcal{U}(0,1)$ in discrete steps of $0.2$ and a binomial distribution $Z_{2}\sim\mathcal{B}(0.7)$. Next, we assign each person a treatment (which will represent the standard of care). This observed treatment allocation differs in Settings 1-3. For a comprehensive description of the data generation process, see Supplementary Material Section 3. Last, we need to simulate the outcome. The generation of the outcome $Y$ relies on the risk function defined by:
@@
  • Setting 2 (S2): In this setting, the clinical utility is zero. Despite employing a different DTR, individual treatments are allocated in a manner that produces an identical counterfactual average outcome to that of the standard of care. This scenario presents a nullifying result.
@@
The case where the DTR under investigation performs worse than the SOC is analogous to Setting 1 and 3 but with a change in sign of the clinical utility. As expected, violations of the positivity assumption in these simulations led to instability or undefined values of the IPW estimators, caused by probabilities near zero in the denominator, however the performance of the correctly specified g-computation estimators remained acceptable as long as the extrapolation is not introducing bias.
@@ 3.1 Results
The results of our simulation study are displayed in Table 1. Notably, for a substantial sample size of $n=2000$, all estimators across Settings 1 to 3 exhibit minimal bias, coupled with a commendable coverage level hovering around $95\%$. However, as sample sizes decrease, we observe a gradual increase in the standard error. The performance degradation of the two IPW estimators is particularly noteworthy, becoming apparent at the smallest sample size of $n=200$, accompanied by a notable increase in bias and standard error, which consequently leads to a decrease in coverage accuracy. With a sample size of $n=200$, we observe that the variance of the $\text{gc}_{b}$ estimator is larger than that of both IPW estimators, which is surprising given our expectation that the IPW estimators would degrade more quickly in general. In our case this can be explained by the $\text{gc}_{b}$ model’s inconsistency to fit all 44 parameters (due to all the interactions) at such small sample sizes. In contrast, the simpler $\text{ipw}_{n}$ model, which requires only 22 parameters, is reliably fitting all parameters. The $\text{gc}_{nb}$ estimator shows the smallest variance because its correctly specified outcome model is exactly the risk function 3 of the data generating mechanism, which makes it impossible to outperform. The estimators $ipw_{b}$ and $ipw_{nb}$ are almost identical in S1-S3. Since $h_{b}$ and $h_{nb}$ are the correctly specified models of the propensity score, they are the same for individuals we sum over, i.e. the ones for which $f(z)=t$. In the misspecified setting $ipw_{b}$ and $ipw_{nb}$ differ as their outcome models for the propensity scores are differently (mis)specified.
@@
[…]
S1 $ipw_{b}$ -0.01 0.14 0.95 0.65 0.29 0.91 2.80 0.45 0.78
$ipw_{nb}$ -0.01 0.14 0.95 0.65 0.29 0.91 2.80 0.45 0.77
$gc_{b}$ -0.02 0.14 0.94 0.05 0.29 0.97 0.31 0.56 0.99
@@ 4.1 Data
The outcome of interest is the log C-reactive protein (CRP) level measured at the earliest time point between 3 months after the treatment switch and up to 1 year after the treatment switch. CRP is a marker of inflammation, hence a smaller outcome is more desirable. Covariates which were thought to possibly influence the outcome and/or the treatment selection included in the models are the phase I treatment, last measured CRP before treatment switch, gender, indicator of being married, indicator of being an immigrant or child of an immigrant, years since RA diagnosis, history of cardiovascular disease, history of chronic obstructive pulmonary disease, and history of osteoperosis. A description of the patients who are included in our emulated target trial is in Table 2.
Table 2: Descriptive statistics of the Danish register data by phase II observed treatment group included in our emulated target trial.
@@ 4.2 Reasoning about time zero
The ideal trial that we would be emulating would have a clearly defined time point where subjects enter the trial and are randomized to either the estimated optimal DTR or the SOC. At the time of entry into the trial, they would also have the necessary biomarkers or other data measured that are inputs to the DTR or SOC regimes. Thus we need to have a similar time zero defined in our emulated trial. The DTR setting differs in an important way from the standard emulated treatment trial because of the requirement of having the necessary biomarkers measured to apply the DTR. If the DTR requires high-dimensional data or a large number of biomarkers, it may be difficult to identify individuals in an observational dataset who have all the required biomarkers measured within a small time window.
@@
In our applied analysis, time zero is clearly defined as the time of treatment switch, since individuals are only eligible for the intervention once. However, in other contexts, such as comparing treatment initiation versus non-initiation, individuals may become eligible multiple times. This complicates the definition of time zero and introduces risks of common biases, such as immortal time bias. One approach to address this challenge is the clone-censor-weight method described by Hernán and Robins [15]. For a broader discussion of strategies for handling time zero, see also [16], [17].
@@ 4.3 Inclusion and exclusion criteria
In the ideal trial, people would be eligible if and when they experience first-line treatment failure, are willing and able to take all possible second-line treatments, and for which a treating clinician does not have a specific treatment indication. Failure is determined based on routine monitoring visits to which we do not have access. However, switching to second-line treatment generally only occurs due to failure. In addition, we use the record of biomarker measurements prior to switching as the indication that clinicians and patients are willing to consider multiple treatments, and need further information on which to base that decision.
To operationalize this ideal trial in our observational data, we include adults who were diagnosed with RA before 1 January 2016 (ICD-10 codes DM05, DM06), and who started first-line treatment with hydroxycloraquine, methotrexate, or sulfasalazine, and who changed or added a biologic or csDMARD during the years 2016, 2017, and 2018. This treatment changing instance would correspond with the phase II treatment decision as described by the EULAR 2023 recommendations [9]. We note that although there may be a few patients that would be a part of our eligible population that are not captured based on our operationalization, we are unable to differentiate them from ineligible subjects due to both groups lacking measurements in our data with no differentiating measure. A flowchart of the exclusion and inclusion criteria can be found in Section 4 of the Supplementary Materials.